OpenStudy (anonymous):

(1-a)2x+(2a+1)x+a+0.5=0

OpenStudy (anonymous):

(1-a)2x+(2a+1)x+a+0.5=0

OpenStudy (maheshmeghwal9):

2x-2ax+2ax+x+a+0.5=0

OpenStudy (maheshmeghwal9):

3x+a+0.5=0

OpenStudy (maheshmeghwal9):

x=(-a-0.5)/3

OpenStudy (maheshmeghwal9):

k!

OpenStudy (anonymous):

distribute then combine like terms to get the value of x

OpenStudy (maheshmeghwal9):

a good theory to apply on such questions which traile told:)

OpenStudy (maheshmeghwal9):

I did the same .I think u can understand now @marzhan

OpenStudy (maheshmeghwal9):

And do it:) good luck

OpenStudy (anonymous):

I don't think that you will be able to get the VALUE of x @traile. There is TWO variables ( "a" & "x") for only ONE equation. You will only end up with a simplified equation 3x + a + 0.5 = 0 x= (-a - 0.5 0)/3 what do you say ?

OpenStudy (anonymous):

@Railey yup, you have a point , but I think that is the only way to get the value of x :)) I guess the letter a will be treated as constant also xD

OpenStudy (anonymous):

@traile : Greek letters generally are used to represent constants. Normal letters (like "a") are used to represent variables. It can be misleading sometimes. Anyway, it's mathematics, so , have fun above all !!

OpenStudy (anonymous):

xD right :))))

OpenStudy (anonymous):

0pening brackets 2x-2xa+2xa+x+a+0.5=0 3x+a+0.5=0 3x+a=0.5 3x=0.5-a x=0.5-a/3

OpenStudy (anonymous):

You forgot the negative sign on "0.5" when you moved it to the other side (2nd & 3rd line)

OpenStudy (anonymous):

sorry was in a hurry

OpenStudy (anonymous):

No worry, it happens.

OpenStudy (anonymous):

@Railey , oh you are new here xD

OpenStudy (anonymous):

Yeap, just happened to end here,... somehow..

OpenStudy (anonymous):

welcome to openstudy XD I'm also new here xD

OpenStudy (anonymous):

Thanks, it's a cool site indeed !